# Markovian Scale Prediction: A New Era of Visual Autoregressive Generation

### v1

@@ 1 Introduction
As a general modeling paradigm, autoregressive modeling has dominated text generation  [wan2023efficient, wan2024d2o, touvron2023llama, chung2024scaling, achiam2023gpt] and multimodal understanding [xie2025collaboration, Chen_2024_CVPR, li2024llava]. Yet, due to suboptimal generation quality, it was not widely adopted in visual generation tasks for a long time. Recently, Visual AutoRegressive modeling (VAR) [VAR] converts next-token prediction to next-scale prediction, further unleashing the potential of the autoregressive modeling to generate high-quality visual content, making autoregression once again a mainstream modeling paradigm for visual generation [kumbong2025hmar, shao2025continuous, ren2024mvardecoupledscalewiseautoregressive, han2025infinity]. However, VAR’s full-context dependency, requiring attention over all previous scales for next-scale prediction, can introduce certain challenges in practice:
@@ 2 Related Work
Early visual autoregressive generation models such as PixelRNN [van2016pixel], PixelCNN [van2016conditional] and PixelCNN++ [salimans2017pixelcnn++] achieved pixel-level conditional modeling, yet this paradigm suffers from limited long dependency and extremely slow sampling. Subsequent works [ramesh2021zero, esser2021imagebart, van2017neural, zhu2024scaling, ding2021cogviewmasteringtexttoimagegeneration, rqvae, yu2022scalingautoregressivemodelscontentrich] treating image synthesis as a token prediction task relied on Vector Quantization (VQ) [van2017neural] to discretize continuous feature maps and autoregressively modeling visual token sequence. However, their raster-scan prediction order conflicts with the spatial structure of the images, leading to inferior performance compared to other alternative-paradigm models [kumari2023multi, guo2024i2v, ddim, ddpm, ho2020denoising, shen2025efficient, goodfellow2014generative]. Recently, VAR [VAR] reformulates next-token prediction as next-scale prediction, further improving the generation quality of visual content. Some VAR-like variants, such as Infinity [han2025infinity], HART [tang2024hart] and VAR-CLIP [zhang2024var] extend VAR to text-to-image generation. Other further improved VAR-like models [li2024controlvar, guo2025fastvar, rajagopalan2025restorevar, xie2024litevar, wu2025nestedautoregressivemodels, jiao2025flexvarflexiblevisualautoregressive, InfinityStar, zhuang2025vargpt] perform various tasks, such as image super-resolution [qu2025visual], 3D Object Generation [chen2025sar3d] and image editing [wang2025training].
@@
The Markov assumption [markov1906rasprostranenie] that posits that the current state depends only on its immediate predecessor, has been widely employed for research in deep learning. Diffusion-based models [croitoru2023diffusion, zhang2025enhancing, zhu2025llada, nie2025large] have achieved remarkable success in various tasks based on this principle. As autoregressive models evolved to involving extremely long sequences, chain-based Markov modeling has been introduced to alleviate the challenges of long-context modeling. MARCOS [liu2025marcos] and MCoT [yang2024markov] introduce structured memory nodes that reformulate chain-of-thought reasoning into a Markovian process for more lightweight inference. In visual autoregressive modeling, although HMAR [kumbong2025hmar] pioneeringly introduces Markov dependency to enhance generation performance, it comes at the cost of increasing the number of inference steps and the token sequence length. In this paper, we revisit this challenging trade-off problem and employ chain-based Markov modeling to develop a visual autoregressive foundation model that reconciles performance with efficiency.
@@ 3.4 Training Strategy
Following the conventional autoregressive paradigm, we construct the ground-truth residual feature sequence $R$ as the supervision target for Markovian scale prediction. Markov-VAR is trained under the teacher-forcing scheme while further adopting the Markovian attention shown in Figure 3, restricting each scale to attend only to its current state. Other main strategies remain consistent with standard visual autoregressive training.
@@ 4.3 Ablation and Analysis
We vary the depth of Markov-VAR from 6 to 24 to analyze the scaling law, corresponding to parameter scales ranging from 19.80M to 1.02B. As shown in Figure 5, charts (a) to (d) show the fitted scaling relation between model size (log10) and performance metrics. Both loss and error rate consistently decrease as model size increases, following obvious power-law trends with high coefficients of determination ($R^{2}>0.99$), indicating that Markov-VAR preliminarily follows the expected scaling law on model size.
@@ 4.4 Visualization
We visualize the Markovian scale prediction process in Figure 6. This visualization provides an intuitive understanding of how Markov-VAR constructs a semantic structure and fills in fine-grained details across multiple scales. The earliest scales capture only very coarse global structures, such as rough color distributions and object silhouettes. As the scale increases, the model gradually refines the semantics. Later scales enrich high-frequency textures and refine local details such as fur patterns and cloud shapes. Notably, the refinement is both smooth and consistent, demonstrating that each Markovian scale effectively preserves essential historical information without relying on full-context dependency.

### v2

@@ 1 Introduction
As a general modeling paradigm, autoregressive modeling has dominated text generation  [wan2023efficient, wan2024d2o, touvron2023llama, chung2024scaling, achiam2023gpt, lu2025easytext, shi2024fonts, shi2025wordcon] and multimodal understanding [xie2025collaboration, Chen_2024_CVPR, li2024llava]. Yet, due to suboptimal generation quality, it was not widely adopted in visual generation tasks for a long time. Recently, Visual AutoRegressive modeling (VAR) [VAR] converts next-token prediction to next-scale prediction, further unleashing the potential of the autoregressive modeling to generate high-quality visual content, making autoregression once again a mainstream modeling paradigm for visual generation [kumbong2025hmar, shao2025continuous, ren2024mvardecoupledscalewiseautoregressive, han2025infinity, hui2025autoregressive, zhang2025easycontrol]. However, VAR’s full-context dependency, requiring attention over all previous scales for next-scale prediction, can introduce certain challenges:
@@ 2 Related Work
Early visual autoregressive generation models such as PixelRNN [van2016pixel], PixelCNN [van2016conditional] and PixelCNN++ [salimans2017pixelcnn++] achieved pixel-level conditional modeling, yet this paradigm suffers from limited long dependency and extremely slow sampling. Subsequent works [ramesh2021zero, esser2021imagebart, van2017neural, zhu2024scaling, ding2021cogviewmasteringtexttoimagegeneration, rqvae, yu2022scalingautoregressivemodelscontentrich, song2025makeanything, wan2024grid] treating image synthesis as a token prediction task relied on Vector Quantization (VQ) [van2017neural] to discretize continuous feature maps and autoregressively modeling visual token sequence. However, their raster-scan prediction order conflicts with the spatial structure of the images, leading to inferior performance compared to other alternative-paradigm models [kumari2023multi, guo2024i2v, ddim, ddpm, ho2020denoising, shen2025efficient, goodfellow2014generative, chen2025transanimate]. Recently, VAR [VAR] reformulates next-token prediction as next-scale prediction, further improving the generation quality of visual content. Some VAR-like variants, such as Infinity [han2025infinity], HART [tang2024hart] and VAR-CLIP [zhang2024var] extend VAR to text-to-image generation. Other further improved VAR-like models [li2024controlvar, guo2025fastvar, rajagopalan2025restorevar, xie2024litevar, wu2025nestedautoregressivemodels, jiao2025flexvarflexiblevisualautoregressive, InfinityStar, zhuang2025vargpt] perform various tasks, such as image super-resolution [qu2025visual], 3D Object Generation [chen2025sar3d] and image editing [wang2025training].
@@
The Markov assumption [markov1906rasprostranenie] that posits that the current state depends only on its immediate predecessor, has been widely employed for research in deep learning. Diffusion-based models [croitoru2023diffusion, zhang2025enhancing, zhu2025llada, nie2025large] have achieved remarkable success in various tasks based on this principle. Chain-based Markov modeling has been introduced to alleviate efficiency challenges of long-context modeling. MARCOS [liu2025marcos] and MCoT [yang2024markov] introduce structured memory nodes that reformulate chain-of-thought reasoning into a Markovian process for more lightweight inference. In visual autoregressive modeling, although HMAR [kumbong2025hmar] pioneeringly introduces Markov dependency to improve generation performance, it comes at the cost of increasing inference steps and the token sequence length. MVAR [zhang2025mvar] innovatively introduces spatial-Markov attention, focusing on reduce computational complexity. We argue that historical information preservation, more concise modeling, and more comprehensive scaling validation are essential for further advancing Markov visual autoregressive modeling. Therefore, we revisit these challenging trade-off problems and employ chain-based Markov modeling to develop a visual autoregressive foundation model that reconciles performance with efficiency.
@@ 3.4 Training Strategy
Following the conventional autoregressive paradigm, we construct the ground-truth residual feature sequence $R$ as the supervision target for Markovian scale prediction.
Markov-VAR is trained under the teacher-forcing scheme while further adopting the Markovian attention shown in Figure 3, restricting each scale to attend only to its current state. Other main strategies remain consistent with standard visual autoregressive training.
@@ 4.3 Ablation and Analysis
We vary the depth of Markov-VAR from 6 to 24 to analyze the scaling law, corresponding to parameter scales ranging from 19.80M to 1.02B. In Figure 5, charts (a) to (d) show the fitted scaling relation between model size (log10) and performance metrics. Both loss and error rate consistently decrease as model size increases, following obvious power-law trends with high coefficients of determination ($R^{2}>0.99$), indicating that Markov-VAR follows the expected scaling law.
@@ 4.4 Visualization
We visualize the Markovian scale prediction process in Figure 6. This visualization provides an intuitive understanding of how Markov-VAR constructs a semantic structure and fills in fine-grained details across multiple scales. The earliest scales capture only very coarse global structures. As the scale increases, the model gradually refines the semantics. Later scales enrich high-frequency textures and refine local details such as fur patterns and cloud shapes. Notably, the refinement is both smooth and consistent, demonstrating that each Markovian scale effectively preserves essential historical information without relying on full-context dependency.